# PertEval: Unveiling Real Knowledge Capacity of LLMs with Knowledge-Invariant Perturbations

### v1

@@ Abstract
Expert-designed close-ended benchmarks serve as vital tools in assessing the knowledge capacity of large language models (LLMs). Despite their widespread use, concerns have mounted regarding their reliability due to limited test scenarios and an unavoidable risk of data contamination. To rectify this, we present PertEval, a toolkit devised for in-depth probing of LLMs’ knowledge capacity through knowledge-invariant perturbations. These perturbations employ human-like restatement techniques to generate on-the-fly test samples from static benchmarks, meticulously retaining knowledge-critical content while altering irrelevant details. Our toolkit further includes a suite of transition analyses that compare performance on raw vs. perturbed test sets to precisely assess LLMs’ genuine knowledge capacity. Six state-of-the-art LLMs are re-evaluated using PertEval. Results reveal significantly inflated performance of the LLMs on raw benchmarks, including an absolute 21% overestimation for GPT-4. Additionally, through a nuanced response pattern analysis, we discover that PertEval retains LLMs’ uncertainty to specious knowledge, potentially being resolved through rote memorization and leading to inflated performance. We also find that the detailed transition analyses by PertEval could illuminate weaknesses in existing LLMs’ knowledge mastery and guide the development of refinement. Given these insights, we posit that PertEval can act as an essential tool that, when applied alongside any close-ended benchmark, unveils the true knowledge capacity of LLMs, marking a significant step toward more trustworthy LLM evaluation.
footnotetext: Work done during Li’s internship at Alibaba Group, under the guidance of Hu.footnotetext: Corresponding author: qiliuql@ustc.edu.cn.
@@ 1 Introduction
Large language models (LLMs) are developing rapidly, and have shown excellent basic capabilities such as reasoning [1, 2], planning [3] and world knowledge [4] in real-world tasks. With the large-scale deployment of LLMs in more and more scenarios, including those safety-critical ones [5], evaluating the real capability of LLMs becomes a necessary and significant task. Knowledge capacity, i.e., the ability to acquire, retrieve and utilize knowledge to solve professional problems, is one of the core capabilities of LLMs. Existing benchmarks for assessing knowledge capacity rely on standardized tests using static, expert-designed datasets [6, 7, 8, 9, 10, 11]. These test datasets consist of multiple-choice questions that include question descriptions, goals, options, and correct answers. They encapsulate valuable domain-specific knowledge that LLMs need to comprehend. The knowledge capacity of LLMs can be directly gauged by their performance on these test datasets.
@@
So far, static data-based benchmarks have successfully evaluated the relative knowledge capacity of LLMs in various domains [12]. However, despite the high quality of expert-designed test data, static data-based benchmarks still face significant challenges in accurately quantifying the true knowledge capacity of LLMs due to their limited test scenarios and the unavoidable risk of data contamination. First, static data-based benchmarks depend on close-ended, single-formed questions for evaluation, which differs a lot to complicated real-world scenarios. For instance, instead of merely selecting an option, some users prefer asking LLMs to judge the correctness of options, while others may directly request the LLMs to generate an answer without providing an option list. Each possible prompt style may affect the performance of LLMs. Furthermore, Schaeffer et al. [13] have demonstrated that using nonlinear or discontinuous evaluation methods could lead to inaccurate evaluation results of LLMs’ capacity. Therefore, to genuinely evaluate the knowledge capacity of LLMs, it is necessary and feasible to test their ability to adjust to a range of complex scenarios that mirror real-world conditions. Second, static data-based benchmarks face an inherent risk of data contamination. These benchmarks publish test data online to ensure transparency and reproducibility in the evaluation process. However, this practice provides opportunities for LLM developers to maliciously fine-tune their models to "memorize" the test data, thereby undermining the benchmark’s integrity. This can result in an overestimation of the LLMs’ knowledge capacity and significantly compromise the reliability of the evaluation results. Efforts to prevent data contamination have primarily focused on detecting such contamination [14, 15] and generating new test data [16]. Nonetheless, high-quality test data contains valuable knowledge that has not been fully leveraged. We contend that it is feasible to utilize this knowledge effectively to assess the true knowledge capacity of LLMs while mitigating the risk of data contamination.
@@
To achieve these goals, we propose PertEval, a knowledge-invariant perturbation-based evaluation toolkit for unveiling real knowledge capacity of LLMs in static data-based benchmarks. This idea stems from an analogy between LLM evaluation and humans’ educational assessment. Similar to LLM evaluation, close-ended questions are prevalent in human evaluation due to their versatility, cost effectiveness and precision of measurement [17], and also suffers from multiple challenges such as cheating [18] and a lack of variety [17]. Solutions include but limited to item personalization [19] and the design of distractive options [20, 21]. Inspired by this, PertEval includes human-like knowledge-invariant perturbations to transform test data to various complicated forms, and transition analysis methods to evaluate and trace the change of the performance of LLMs in different test scenarios. Knowledge-invariant perturbations consist of the rewriter LLM-based content-level perturbation for largely changing text features of questions to prevent data contamination, and a series of format-level perturbations for comprehensively covering possible test scenarios in real-world conditions. Transition analysis methods consist of not only evaluation metrics for measuring real knowledge capacity of LLMs, but also response pattern analysis for tracing the cause of the change of LLM’s performance, which probes in-depth weaknesses of the knowledge capacity of LLMs and provide insights for the refinement of LLMs.
@@
In experiments, we re-evaluate six state-of-the-art LLMs using PertEval. Transition analysis results first reveal a significant overestimation of the knowledge capacity of LLMs in MMLU, with an absolute 21% overestimation for GPT-4 and 35% overestimation for Gemini-1.0-Pro. Response pattern analyses further unveil that the overestimation for most LLMs is caused by a significant increment of the ratio of extra incorrect choices in response patterns, with an increment up to 9.7% for GPT-4 and 27.7% for Gemini-1.0-Pro. These results illustrate the potential rote memorization of the LLMs on correct options. Detailed transition analyses in terms of overall performance stability and correct response consistency reveal various weaknesses in the knowledge capacity of existing LLMs, providing guidance for their refinement. Given these insights, we posit that PertEval serves as an essential tool for evaluating the real knowledge capacity of LLMs when applying to any existing close-ended benchmarks, marking a significant step towards more trustworthy LLM evaluation.
@@ 2 Related Work
Benchmarks for Knowledge Capacity Evaluation. Benchmarks for knowledge capacity evaluation could be classified into general knowledge benchmarks or professional knowledge benchmarks [12]. General knowledge benchmarks aim to evaluate the general knowledge capacity of LLMs in all kinds of domains, such as MMLU [6], C-Eval[22] and ARC[8]. On the other hand, professional knowledge benchmarks aim to deeply evaluate the professional capacity of LLMs in specific domains, such as MedMCQA [9] (Medicine), ScienceQA [23] (Science) and GSM-8K [11] (Mathematics). These benchmarks depend on professional multiple-choice questions to measure the capacity of LLMs.
Trustworthiness of Knowledge Capacity Evaluation. Research about the trustworthiness of knowledge capacity evaluation are mainly from the test data acquisition perspective and the response generation perspective. For test data acquisition, the data contamination risk have been proved widely exist in test data [24, 25, 26]. Recent works emphasize the detection of contamination [14, 15] and generating contamination-free test data [16, 27]. For response generation, it has been observed that multiple text-related biases could affect LLMs’ performance [28, 29, 30, 31], such as option ordering bias [28] and selection bias [29]. Debiasing methods like PriDe [14] serve as external plug-in for LLMs in multiple-choice question answering, which are inappropriate for evaluating real knowledge capacity.
Adversarial Text Attack for LLMs. Adversarial text attacks aim to mislead language models to generate wrong outputs. Important research directions of them in the era of LLM include but not limited to jailbreak attack and text classification attack. Traditional text attacks could be classified into character-level [32, 33], word-level [34] and sentence level [35]. Recently, LLM-based perturbations such as PromptAttack [36] have also been introduced. However, considering their cost and the comprehensiveness, existing text attack methods are insufficient for achieving our research goal.
@@ 3 Methodology
Table 1: Standards of knowledge invariance
@@
3.1 Knowledge-invariant Perturbation
3.1.1 Content-level Perturbation: Knowledge-invariant Paraphrasing
The content-level perturbation aim to largely change the statement of questions while keeping the original knowledge to prevent data contamination of orginal test data. The key challenge of knowledge-invariant paraphrasing is how to preserve the original knowledge while change the statement as possible it can. To this end, we first propose standards of knowledge invariance, as presented in Table 1 for the checking of knowledge invariance. Next, we propose a sentence-by-sentence paraphrasing algorithm using an LLM rewriter. The motivation is that the logic of a question text is the composition of the logic of the sentence sequences. Formally, let $q=(s_{1},s_{2},\ldots,s_{T})$ be the question text, where $s_{t},t=1,2,\ldots,T$ denotes the $t$-th sentence. Then the semantic of $s_{t}(t\geq 2)$ depends on its prerequisite sequence $(s_{1},\ldots,s_{t-1})$. To rewrite the whole question text, an LLM rewriter is instructed to rewrite sentence by sentence given each sentence and its original context. The paraphrasing algorithm, the prompt template for the rewriter LLM and an example of the perturbation are available at Appendix C.1.
3.1.2 Format-level Perturbation: Question Format Refactoring
Question format refactoring aim to evaluate the robustness of LLMs’ knowledge capacity when confronting complicated test conditions. To this end, we design a variety of format-level perturbation strategies that cover most possible format bias of LLMs in real-world scenarios.
Option Permutation (OptionPerm). OptionPerm reorders option contents while keeping the order of option IDs. It aims to evaluate the option ordering bias in the knowledge acquirement of LLMs. The default setting of OptionPerm is to reverse the order of option contents to completely destroy the local order of contents. When applying OptionPerm, the true answer of the question is also changed to be aligned with the perturbation result.
Option Format Refactoring (OptionForm). OptionForm changes the format of option IDs (e.g., append a right parenthesis to the end of IDs). It aims to evaluate the option format bias in the knowledge acquirement of LLMs. OptionForm might affect the performance of LLMs by changing the dependency beteween different tokens in options.
Option ID Shifting (OptionCaesar). OptionCaesar shifts the ASCII value of option IDs to change their character, which is similar to Caesar encryption. It aims to evaluate the selection bias in the knowledge acquirement of LLMs, whose existence has been empirically proved in some LLMs. By applying OptionCaesar, we replace common option IDs (e.g., A/B/C/D) by uncommon option IDs so that we can observe whether the value of IDs affect the performance of LLMs.
Question Type Changing (ChangeType). ChangeType changes a multiple choice question to a multiple judgement question. It aims to evaluate the question type bias of LLMs. Since the feasible solution space of a multiple choice question and the corresponding multiple judgement question is identical (given $N$ options, the size of the feasible solution space is $(2^{N}-1)$), an LLM that robustly acquire knowledge/skills in a question should be insensitive to ChangeType. Therefore, it is a significant perturbation for LLMs’ knowledge acquirement evaluation.
Question Position Swapping (SwapPos) SwapPos swaps the position of question texts with options. It aims to evaluate the global ordering bias of LLMs. For rational human test-takers, SwapPos does not change their performance on questions because merely move question text to after the options does not change semantic information and knowledge of the question. However, although token-level knowledge is also kept in SwapPos, this perturbation still might be a difficult challenge for LLMs, because it destroys the global ordering of input prompts, which might largely change the distribution of output of self-regression text generation models.
3.2 Knowledge Invariance Checking
@@
LLM-based Knowledge Invariance Scoring. This method checks the knowledge invariance of perturbations from the perspective of LLM’s perception. Specifically, we first choose a superior LLM (e.g., gpt-4-turbo) as the referee. Then, give the test dataset $D_{dual}=\{(q_{i},q_{i}^{\prime})|i=1,2,\ldots,|D|\}$, we construct the scoring prompt based on the prompt template, standards of knowledge invariant perturbations (see Table 1) and the criteria of scores for knowledge invariance judgement (see Table 8 in Appendix C.1) for each pair of original and perturbed questions. After collecting the output score of the referee LLM from the scoring prompt, the average score is calculated as the final result.
@@
Testing on Mastered Questions for LLMs. Different from the scoring method that is input-oriented, this method evaluates the output-oriented knowledge invariance. The motivation is that if a perturbation is knowledge-invariant, an LLM’s performance on questions whose knowledge it really mastered should be the same on the original and the perturbed version. The challenge is that mastered questions for LLMs are hard to find. To this end, we propose a substitute solution that utilizes questions that most LLMs can correctly answer to represent the set of mastered questions for an LLM. Then an LLM test-taker (gpt-4-turbo in experiments) is required to answer the original and perturbed versions of these questions to validate the knowledge invariance of the perturbation. In a word, the test-based knowledge invariance checking procedure can be summarized as follows:
@@
    Given a set of LLM test-takers and the evaluation dataset $D$, collect question samples that all LLMs answer correctly to construct the mastered question set $D_{simple}$.
@@
    Apply a perturbation to each of sample of $D_{simple}$ to obtain the perturbed dataset $D_{simple}^{\prime}$.
@@
    For each LLM, evaluate its performance on $D_{simple}^{\prime}$. If there exists at least one LLM whose performance on $D_{simple}^{\prime}$ has no significant difference with that on $D_{simple}$, then the perturbation is knowledge-invariant.
This method focuses more on the influence of perturbation on the text generation process of LLMs. Compared to knowledge invariance scoring operations, this methods is more suitable for evaluating the knowledge invariance of content-level perturbations.
3.3 Transition Analysis For Measuring Real Knowledge Capacity
The transition analysis aims to measure the real knowledge capacity of LLMs from their responses to benchmark datasets. In this part, we first propose the metric of real knowledge capacity. Next, we propose two aspects for detailed analyses of LLMs’ weaknesses in knowledge capacity.
Metric of Real Knowledge Capacity. To measure real knowledge capacity of LLMs, we propose the Consistent Accuracy (ACC@Consist) as the evaluation metric. The motivation is that if a question and its corresponding knowledge is truly mastered by an LLM, then the LLM’s performance should be consistent on any perturbed version of the question, including the original version and the most complicated version. To this end, let $M(\cdot)$ denotes the response function of an LLM. Let $x=(q_{x},y_{x})\in X$ denotes a test question, where $q_{x}$ denotes the question text and $y_{x}$ denotes the correct answer(s). Let $\sigma^{*}:X\to X$ denotes the most complicated composite knowledge-invariant perturbation. Let $D=\{x_{1},x_{2},\ldots,x_{n}\}$ denotes the test set. Then ACC@Consist is defined as:

### v2

@@ Abstract
Expert-designed close-ended benchmarks are indispensable in assessing the knowledge capacity of large language models (LLMs). Despite their widespread use, concerns have mounted regarding their reliability due to limited test scenarios and an unavoidable risk of data contamination. To rectify this, we present PertEval, a toolkit devised for in-depth probing of LLMs’ knowledge capacity through knowledge-invariant perturbations. These perturbations employ human-like restatement techniques to generate on-the-fly test samples from static benchmarks, meticulously retaining knowledge-critical content while altering irrelevant details. Our toolkit further includes a suite of response consistency analyses that compare performance on raw vs. perturbed test sets to precisely assess LLMs’ genuine knowledge capacity. Six representative LLMs are re-evaluated using PertEval. Results reveal significantly inflated performance of the LLMs on raw benchmarks, including an absolute 25.8% overestimation for GPT-4. Additionally, through a nuanced response pattern analysis, we discover that PertEval retains LLMs’ uncertainty to specious knowledge, and reveals their potential rote memorization to correct options which leads to overestimated performance. We also find that the detailed response consistency analyses by PertEval could illuminate various weaknesses in existing LLMs’ knowledge mastery and guide the development of refinement. Our findings provide insights for advancing more robust and genuinely knowledgeable LLMs. Our code is available at https://github.com/aigc-apps/PertEval.
footnotetext: Work done during Li’s internship at Alibaba Cloud Computing, under the guidance of Hu.footnotetext: Qi Liu is the corresponding author.
@@ 1 Introduction
Figure 1: An overview of the PertEval evaluation toolkit. PertEval uses content-level and format-level perturbations to generate perturbed dataset $D^{\prime}$ from existing close-ended benchmark dataset $D$. Next, it evaluates the knowledge capacity of LLMs via response consistency analysis. PertEval also demonstrates in-depth the performance feature of LLMs via response pattern analysis.
@@
Large language models (LLMs) are developing rapidly, and have shown excellent basic capabilities, such as reasoning [1, 2], planning [3] and world knowledge [4], in real-world tasks. Given the widespread deployment of LLMs across an increasing number of scenarios, including those that are safety-critical [5], evaluating the real capability of LLMs becomes a necessary and significant task. Knowledge capacity, i.e., the ability to retrieve and utilize acquired knowledge to solve professional problems, is one of the core capabilities of LLMs [4]. Existing knowledge capacity evaluation largely rely on standardized tests using close-ended benchmarks [6, 7, 8, 9, 10, 11]. These benchmarks consist of multiple-choice questions that include question descriptions, goals, options, and correct answers. They encapsulate valuable domain-specific knowledge that LLMs need to comprehend. The knowledge capacity of LLMs can then be directly gauged by the performance on these test datasets.
@@
To date, static benchmarks have been essential in assessing the relative knowledge capacity of LLMs [12] because of their high quality and cost effectiveness. However, due to limited test scenarios [13] and the unavoidable risk of data contamination [14], these benchmarks still encounter significant challenges in accurately measuring the true knowledge capacity of LLMs. First, static benchmarks rely exclusively on close-ended questions in fixed formats for evaluation, which differs a lot from real-world scenarios [15]. For instance, instead of merely selecting an option, some users prefer asking LLMs to judge the correctness of options, while others may directly request the LLMs to generate an answer without providing an option list [16]. These varying prompting styles have an innegligible effect on performance [17]. Furthermore, previous work [18] has demonstrated that nonlinear or discontinuous evaluation metrics could lead to inaccurate implications about LLM capacity. Therefore, to genuinely evaluate the knowledge capacity of LLMs, it is necessary to test their performance across a variety of scenarios that mirror real-world conditions. Second, in terms of data contamination [14], knowledge capacity benchmarks usually publish their test data online to ensure transparency and reproducibility (e.g., [6, 7]). However, this practice enables LLMs to memorize the test data during pre-training and alignment, which can lead to an overestimation of knowledge capacity, thus undermining the trustworthiness of evaluation results [19]. Efforts to mitigate data contamination have primarily focused on detecting such contamination [14, 20] and generating new test data [21]. Nonetheless, high-quality benchmark data contain rich knowledge that is valuable for evaluation. We suppose that it is feasible to utilize this knowledge effectively to assess the true knowledge capacity of LLMs while reducing the risk of data contamination.
@@
To achieve this goal, we introduce PertEval, an evaluation toolkit that utilizes knowledge-invariant perturbations on static benchmarks to unveil the real knowledge capacity of LLMs, as illustrated in Figure 1. This idea stems from an analogy with human educational assessment. Like LLM evaluation, close-ended questions are prevalent in human assessments due to their versatility, cost-effectiveness and precision of measurement [22]. However, they also face challenges such as cheating [23] and a lack of variety [22]. Solutions to these issues include item personalization [24] and the creation of distractive options [25, 26]. Inspired by this, PertEval incorporates human-like knowledge-invariant perturbations to restate static test data to various forms and employs response consistency analyses to evaluate and trace the change of LLMs’ performance in different test scenarios. Specifically, knowledge-invariant perturbations consist of both the content-level perturbation that extensively writes questions to minimize data contamination and a series of format-level perturbations that comprehensively cover potential real-world test scenarios. Response consistency analyses encompass basic evaluation metrics to measure the real knowledge capacity of LLMs and response pattern analysis to investigate the causes behind changes in LLM performance. Consequently, PertEval could deeply probe the LLMs’ weaknesses in knowledge mastery and offer insights for their refinement.
@@
In experiments, we re-evaluate six representative LLMs using PertEval. Capacity metric results first reveal a significant overestimation of the knowledge capacity of LLMs by static benchmarks, with an absolute 25.8% overestimation for GPT-4 and 38.8% for Gemini-1.0-Pro on MMLU [6] (results of Gemini-1.0-Pro are available at Appendix D.2). Response pattern analyses further unveil that the performance decline with PertEval is mainly caused by an increase of selecting extra incorrect choices, with an increase of up to 9.7% and 27.7% for GPT-4 and Gemini-1.0-Pro, respectively. These findings highlight the potential for rote memorization of correct options by LLMs on static benchmarks. Detailed response consistency analyses in terms of overall performance stability and correct response consistency indicate various weaknesses in the knowledge capacity of existing LLMs, such as the vulnerability to content-level perturbation and the change of global text order.
Overall, our contributions are as follows:
  • We propose an evaluation toolkit that utilizes knowledge-invariant perturbations on close-ended evaluation benchmarks to unveil the real knowledge capacity of LLMs, a significant step towards more trustworthy LLM evaluation.
  • We re-evaluate the knowledge capacity of six representative LLMs using PertEval. Evaluation results not only demonstrate significantly inflated performance of LLMs by static benchmarks, but also reveal LLMs’ uncertainty to specious knowledge and rote memorization to correct options.
  • We demonstrate the vulnerability of various LLMs to different perturbation strategies in PertEval and provide insights for the refinement in terms of promoting LLMs’ knowledge capacity.
@@ 2 Related Work
Knowledge Capacity Evaluation of LLMs Research works of the knowledge capacity evaluation LLMs consist of two lines, i.e., evaluation benchmarks and evaluation methodologies. The first line aim to design comprehensive and accurate benchmarks to quantize the knowledge capacity of LLMs. Evaluation benchmarks could be further classified into general or professional knowledge benchmarks [12]. The first category aims to evaluate the general knowledge capacity of LLMs across a wide range of domains, such as MMLU [6], C-Eval [27] and ARC [8]. The second category aims to deeply evaluate the professional capacity of LLMs in specific domains, such as MedMCQA [9] in medicine and ScienceQA [28] in science. These benchmarks depend on professional multiple-choice questions to quantitatively measure the capacity of LLMs. The other research line, i.e., evaluation methodologies, aim to improve the accuracy and truthfulness of evaluation via data-driven or model-driven techniques. Along this line, recent studies have emphasized data-contamination detection techniques [14, 20] and contamination-free test data generation [21, 29]. In addition, psychometric-based techniques [30], dynamic evaluation techniques [31, 32], and perturbation-based evaluation techniques such as CheckList [33] and PolyJuice [34] have been proposed to comprehensively evaluate the capability of language models (LMs) from different aspects. However, these methodologies are unsuitable for knowledge capacity evaluation, and are hard to utilize valuable information in expert-designed datasets for trustworthy evaluation.
Consistency of LMs The consistency of an LM denotes its behavior invariance under meaning-preserving alternations in its input [35], which is significant in natural language understanding. Since the emergence of pretrained language models (PLMs), massive efforts have been made to evaluate the consistency of LMs on various fields. Elazar et al. [35] revealed the poor consistency performance of PLMs on factual knowledge. Fierro and Søgaard [36] extended the study to multilingual PLMs and obtained similar findings. Jang et al. [37] proposed a taxonomy for various concepts of consistency and established a benchmark, BECEL, to evaluate the consistency of PLMs. As LLMs develop swiftly, Wang et al. [38] proposed to use self-consistency decoding to empower the Chain-of-Thought reasoning ability of LLMs. Recently, Rajan et al. [39] proposed KonTest, an autonomous evaluation framework that utilizes knowledge graphs to generate test samples, to probe the inconsistency in LLMs’ knowledge of the world. In summary, these research works view consistency as a research subject that needs to be measured or intervened. Differently, in our study, the consistency of LLMs is viewed as a technical method, i.e., a measurement for probing LLMs’ real knowledge capacity.
Adversarial Text Attacks on LMs These techniques aim to mislead language models to yield wrong or toxic outputs with small text perturbations, which play an indispensable role in the research of robustness and safety of language models. Such efforts include but not limited to jailbreaking [40] and text classification attack [41]. Text attacks could also be classified into character-level [42, 43], word-level [44] and sentence-level [45]. Recently, LLM-based perturbations like PromptAttack [41] have also been introduced. However, considering their cost and the comprehensiveness, existing text attack methods are insufficient for the purpose of this work. Moreover, our PertEval is built upon a new concept of knowledge-invariant perturbation, which is an important supplement to this topic.
@@ 3 Methodology
3.1 Knowledge-invariant Perturbations
We first present the two types of knowledge-invariant perturbations to restate static test questions. The knowledge-invariant property of these perturbations will be discussed in the next subsection.
Content-level perturbation: knowledge-invariant paraphrasing. The goal of content-level perturbation is to substantially alter the phrasing of questions while retaining the original knowledge, thereby mitigating data contamination in the test data. The key challenge of such knowledge-invariant paraphrasing is to preserve the original knowledge while changing the statement as much as possible. A test question is a composite of sentences, with each provide either backgrounds, conditions or goals of the question. Therefore, to preserve knowledge of the original question, we propose a sentence-by-sentence paraphrasing algorithm using an LLM rewriter. Formally, let $q=(s_{1},s_{2},\ldots,s_{T})$ represent the question text, where $s_{t}$ denotes the $t$-th sentence (for $t=1,2,\ldots,T$). The semantic of $s_{t}$ (for $t\geq 2$) depends on its prerequisite sequence $(s_{1},\ldots,s_{t-1})$. To rewrite the entire question, an LLM rewriter is instructed to paraphrase sentence by sentence given each sentence and its original prerequisite sequences. Detailed descriptions of the paraphrasing algorithm, the prompt template for the rewriter LLM, and an example of the perturbation can be found in Appendix B.1.
Format-level perturbation: question format refactoring. The objective of question format refactoring is to assess the robustness of LLMs’ knowledge capacity under complicated test conditions. To achieve this, we have developed a variety of format-level perturbation strategies to comprehensively evaluate the resilience of LLMs’ knowledge capacity in various test scenarios.
  • Option permutation (OptionPerm). OptionPerm reorders the contents of options while maintaining the original order of option IDs. Its goal is to evaluate option ordering bias in the knowledge acquisition of LLMs. By default, OptionPerm reverses the order of option contents to completely disrupt their local sequence.
  • Option format refactoring (OptionForm). OptionForm modifies the format of option IDs, such as by appending a right parenthesis to the end. This perturbation aims to assess the option format bias in the knowledge acquisition of LLMs. OptionForm may influence LLM performance by altering the dependency between different tokens within the options.
  • Option ID shifting (OptionCaesar). OptionCaesar shifts the ASCII value of option IDs to change their character, which is similar to Caesar encryption. This technique aims to investigate the selection bias in the knowledge acquirement of LLMs, a phenomenon empirically observed in some models. By replacing common option IDs with less common ones (e.g., A/B/C/D $\to$ U/V/W/X), OptionCaesar allows us to observe whether the values of the IDs impact the LLMs’ performance.
  • Question type changing (ChangeType). ChangeType converts a multi-choice question into a multi-judgment question. This perturbation aims to examine the question type bias in LLMs. Since the feasible solution space of a multi-choice question and the corresponding multi-judgement question is identical (given $N$ options, the size of the feasible solution space is $12^{N}-$ 1), an LLM that robustly acquire knowledge/skills in a question should be insensitive to ChangeType.
  • Question position swapping (SwapPos). SwapPos switches the position of the question text and the options. It aims to evaluate the global ordering bias of LLMs. For rational human test-takers, SwapPos does not affect performance, as it does not alter the question content. However, SwapPos can significantly change the output distribution of self-regressive text generation models by disrupting the global ordering of input prompts.
Examples of format-level perturbations are available at Appendix B.2.
3.2 Knowledge Invariance Verification
Table 1: Standards of knowledge-invariance scoring.
@@ 3.2 Knowledge Invariance Verification
We next investigate the knowledge-invariance property of the proposed perturbations.
@@
Knowledge invariance scoring. This approach checks the knowledge invariance of perturbations from the perspective of humans’ and LLMs’ perception. Specifically, based on previous works in the measurement of semantic similarity [46] and characteristics of close-ended benchmarks, we first propose standards of knowledge-invariance scoring, as shown in Table 1. Next, we recruit professional human volunteers and utilize superior LLMs, such as claude-3-sonnet, respectively, to serve as the referee to rate knowledge invariance scores. Given a set of original and perturbed question pairs $D_{dual}=\{(q_{i},q_{i}^{\prime})\mid i=1,2,\ldots,|D|\}$, we construct scoring prompts based on a predefined template, knowledge invariant standards, and the scoring criteria for knowledge invariance judgement (see Table 9 in Appendix B.1). After collecting the output scores from the referee LLM using these scoring prompts, we calculate the average score as the final result.
@@
Testing on mastered questions for LLMs. This approach evaluates knowledge invariance based on the output performance of LLMs. The rationale here is that if a perturbation is knowledge-invariant, an LLM’s performance on questions that it has truly mastered should remain consistent between the original and perturbed versions. Identifying mastered questions for LLMs, however, poses a challenge. To this end, we propose using questions that most LLMs can correctly answer as a proxy for mastered questions. An LLM, such as gpt-4-turbo, is then required to answer both the original and perturbed versions of these questions to validate the knowledge invariance of the perturbation. In summary, the test-based knowledge invariance checking procedure involves the following steps:
@@
    Given a set of LLM test-takers and the evaluation dataset $D$, construct the mastered question set $D_{simple}\subset D$ consisting of questions that all LLMs could answer correctly.
@@
    Apply a perturbation to each sample in $D_{simple}$ to create the perturbed dataset $D_{simple}^{\prime}$.
@@
    Assess the performance of a knowledgeable LLM on both $D_{simple}$ and $D_{simple}^{\prime}$. If the test results show no significant difference in performance, the perturbation can be considered knowledge-invariant.
This method emphasizes the impact of perturbations on the question-answer process of LLMs. In comparison to the scoring test, this approach is more appropriate for evaluating the knowledge invariance of content-level perturbations. We present detailed results of both tests in Experiments.
3.3 Response Consistency Analyses for Measuring Real Knowledge Capacity
We further devise a suite of response consistency analyses that, by comparing performance on raw vs. perturbed test sets, unveils the real knowledge capacity of LLMs. These include a metric for quantifying calibrated knowledge capacity and the more fine-grained ones helpful for revealing the vulnerability of LLMs to different perturbation strategies.
Metric of real knowledge capacity. To measure real knowledge capacity of LLMs, we propose the Consistent Accuracy (ACC@Consist) as the evaluation metric. The rationale is that if an LLM has truly mastered a question and its underlying knowledge, its performance should remain consistent across all versions of the question, including both the original and the most complex perturbed versions. Formally, let $M(\cdot)$ denotes the response function of an LLM. Let $x=(q_{x},y_{x})\in X$ represent a test question, where $q_{x}$ and $y_{x}$ denote the question text and the correct answer(s) of $x$. Further let $\sigma^{*}:X\to X$ be the most complicated composite knowledge-invariant perturbation and $D=\{x_{1},x_{2},\ldots,x_{|D|}\}$ the raw test set. Then ACC@Consist is defined as: